# Can LLMs Understand the Implication of Emphasized Sentences in Dialogue?

### v1

@@ Abstract
Guan-Ting Lin, Hung-yi Lee National Taiwan University, Taiwan {f10942104, hungyilee}@ntu.edu.tw
@@ 1 Introduction
Large Language Models (LLMs)(Wei et al., 2022; Touvron et al., 2023a; OpenAI, 2023) have revolutionized natural language processing by using extensive unlabeled text data. They generate and understand natural language accurately, capturing nuances and complex relationships, and are widely used in applications such as dialogue systems and virtual assistants. Despite their capabilities, the ability of LLMs to grasp the subtle nuances and complex meanings conveyed through “emphasis" in human dialogue remains unclear. Emphasis is a critical component in human dialogue, used to highlight important information and intentions. For instance, altering the emphasis in the sentence “I never said he stole my bag" from “he" to “my" can drastically change its meaning. Understanding such nuances is essential for LLMs to effectively communicate with humans.
@@
To address this gap, we introduce a novel benchmark evaluation dataset, Emphasized-Talk, featuring real dialogue samples with annotations that capture the implications of emphasis interpreted by humans. Emphasized-Talk contains the same dialogue context and the same current turn sentence in different emphasized words or phrases. This dataset is designed to test whether the LLMs can accurately understand the meaning and intention of emphasis in dialogue. Our study includes both open-sourced and commercial LLMs of varying sizes to provide a comprehensive analysis of current capabilities and limitations.
Furthermore, we propose an automatic evaluation method that utilizes the capabilities of GPT-4 OpenAI (2023) to assess the performance of these models. By comparing various automatic evaluation methods, we find that GPT-4’s evaluations correlate well with human evaluation scores. Our findings indicate that open-source LLMs generally struggle to understand the implications of emphasized sentences. While commercial LLMs perform better, there remains room for improvement in their comprehension of emphasized text.
@@
    We introduce an automated evaluation pipeline using GPT-4, enabling efficient and consistent assessment of models on their handling of emphasis, which significantly reduces the need for manual testing.
@@ 2 Related works
Human Communication Beyond Lexical Information: In addition to text, the extra tags like emotion (Pohl et al., 2017; Xue et al., 2023), emoji (Demszky et al., 2020), sentiment (Lin et al., 2024b), prosody (Ward et al., 2012; Lin et al., 2024a), and non-lexical sounds (Ward, 2006) are critical in understanding and generating human dialogue. These non-verbal cues provide essential context and pragmatic use to the interaction, often conveying nuances that textual information alone cannot fully capture (Ward, 2004). For instance, emotions and sentiment can indicate the speaker’s attitude and feelings, while speaking style and prosody can reveal additional layers of meaning, such as sarcasm, emphasis, or urgency. This work focuses on emphasis as the non-lexical cue in human communication, aiming to understand the implication meaning of beyond text.
Emphasis in Dialogue: In communication, whether spoken or written, emphasis is a crucial tool for enhancing meaning, indicating emotional states, or highlighting important dialogue segments. In dialogue systems, understanding the subtleties conveyed through emphatic cues is essential for achieving truly natural interactions (Pierrehumbert and Hirschberg, 1990; Wagner and Watson, 2010; Jackson, 2016). Emphasis aids in disambiguating syntactic structures, reinforcing intentions, and expressing nuances that are vital for comprehensive understanding in conversational contexts (Buchanan, 2013). Due to the importance of emphasis, studies on emphasis detection (Arons, 1994; Suni et al., 2017; Talman et al., 2019; Zhang et al., 2018; Vaidya et al., 2022; Morrison et al., 2024), emphatic Text-to-Speech synthesis (Fernandez and Ramabhadran, 2007; Seshadri et al., 2022; Stephenson et al., 2022), and Speech-to-Speech translation (Goldman et al., 2016; de Seyssel et al., 2023) are long-standing research topics to capture emphasis. This work focus on evaluating the LLMs’ ability to understand the high-level meaning and intention of emphasis.
@@ 3.1 Sentence Emphasis Selection
For each dialogue, we select the current sentence to be emphasized based on the following criteria: (1) Sufficient dialogue context: The current turn is selected only if there are more than two preceding dialogue turns, ensuring adequate contextual information. (2) Availability of emphasized targets: Since emphasizing function words like punctuation, articles, and proper nouns rarely affects sentence meaning, we select sentences containing more than four non-functional words and proper nouns for further consideration.
@@ 3.2 Human Annotation with Implication
Selecting which words and phrases to emphasize is non-trivial. In human communication, emphasis placement depends on the dialogue context and the information the speaker wishes to convey. In this work, human annotators choose where to place emphasis and document the implied meaning behind their choices, making the data more pragmatic and reflective of real-world dialogue. Each annotator selects two different words or phrases to emphasize and note the implied meaning. The task template and instructions are shown in Appendix Figure 3.
@@ 3.3 Filtering by Agreement
When multiple annotators select the same word or phrase to emphasize, there are often several annotations of implied meanings. If the implied meanings are significantly different between annotators, indicating a lack of consensus or too many possible implied meanings based on annotators’ backgrounds, we check the semantic similarity among the annotated meanings using GPT-4 (see Appendix). Samples without agreement are filtered out. The percentage of samples that passed the filtering is 79.1%. Overall, we collect 984 dialogue sample pairs.
@@
Table 2: Qualitative example with Emphasizing on different targets in the current turns. “0" and “1" denote the speaker’s identity. For implication rows, the word or phrase in quotation refers to the emphasized target, and the sentence after colon is the implication. The “Predicted implication" is from the Claude 3 Sonnet model.
@@ 4 Evaluation
For human evaluation, we recruit human annotators to rate the scoring from 0 to 5, given the rating guideline (the detailed instruction is shown in Appendix Figure 4). Each sample is rated by three annotators, the average score is the Mean Opinion Score (MOS). However, human evaluation incurs high costs and significant time requirements, requiring skilled evaluators who can understand and analyze the subtle nuances of language and context within the dialogues. Therefore, we propose using the state-of-the-art LLM, GPT-4 (OpenAI, 2023), for automatic evaluation. The automatic evaluation methods are illustrated in Figure . We adopt a similar LLM automatic evaluation method as in previous works (Chiang and Lee, 2023a, b), instructing the LLM to predict ratings based on guidelines. This is denoted as auto-gpt4 score. Given potential uncertainties in the LLM’s comprehension of emphasized sentences, we also use a similarity score compared to ground truth annotations, employing both the BERT score (BERTf1) and GPT-4 (auto-gpt4-gt). Appendix F describes more details on automatic evaluation.
@@ 5 Results
5.1 Large Language Models to be Evaluated
We use both open-source and closed commercial LLMs for evaluation. Specifically, we evaluate different versions of Llama (Touvron et al., 2023a) (Llama 2 (Touvron et al., 2023b) and Llama 3 (AI@Meta, 2024)) and various model sizes (ranging from 7B to 70B parameters). In addition to Llama, we experiment with Mistral 7B (Jiang et al., 2023). All models used are the instruction-tuned chat versions, not the pre-trained models. For commercial LLMs, we evaluate Claude 3 Sonnet [^2] https://www.anthropic.com/news/claude-3-family and ChatGPT 3.5 (gpt-3.5-turbo-0125) [^3] https://openai.com/index/chatgpt/. For all the LLMs, we provide the prompt as detailed in the Appendix D section.
@@ 5.2 LLMs’ Performance on Implication Modeling
Table reports show the scores of different LLMs. Commercial LLMs generally outperform open-source ones in MOS. The top model is Claude 3 Sonnet with a MOS of 3.73, followed by ChatGPT at 3.59. Among open-source LLMs, Llama 3-70B scores 3.51, close to ChatGPT. The Llama 3 series consistently outperforms the Llama 2 series of similar sizes. Mistral 7B ranks between Llama 2-7B and Llama 3-8B, even surpassing Llama 2-70B. None of the LLMs reach the optimal score of 5, with the best at 3.73, indicating room for improvement in understanding emphasized sentences. More error analyses are shown in Appendix B.
@@ 5.4 Relative Ranking for Automatic Evaluation Methods
The ranking order of LLMs by automatic evaluation scores aligns closely with human scores. Spearman’s rank correlation between human MOS and three automatic methods shows high coefficients: 0.964 for auto-gpt4 and BERTf1, and 0.857 for auto-gpt4-gt. Auto-gpt4 and BERTf1 had only two incorrect model rankings. Despite low Pearson’s and Kendall’s correlations, BERTf1’s rankings are similar to human MOS. Overall, auto-gpt4 achieves the both highest score and rank correlation.
@@ 5.5 Qualitative Example
Table 2 shows examples of the same dialogue context with different emphasized words or phrases. For "single," the Claude 3 Sonnet model incorrectly interprets it as referring to a cart for individual use, whereas it actually refers to the number of carts. For "right now," the model correctly understands it implies immediate use. This demonstrates that while commercial LLMs can infer different meanings based on emphasis, their accuracy depends on the complexity and depth of the context.
@@ Acknowledgement
We thank Cheng-Han Chiang for his insightful discussion. Guan-Ting Lin is supported by the NTU GICE and the NTU Ph.D. scholarship program.
Appendix A Example Appendix
@@
Appendix B Analysis on Unsatisfied Model Prediction
We analyze the cases resulting in unsatisfactory ratings (average MOS below 2). The examples are shown in Table 4. We observe that Claude 3 Sonnet overinterprets the emphasized information in an extremely contrasting way, but this is not relevant to the dialogue context. In contrast, the Llama 2-7B and Mistral-7B models fail to focus on the current turn and the emphasized word, merely summarizing the entire dialogue. These results indicate that small open-sourced LLMs struggle to understand emphasized text and its implications, and sometimes even fail to follow instructions. On the other hand, commercial LLMs can follow instructions but may overinterpret the emphasized text. The ideal interpretation should be reasonable and relevant to the dialogue context.
Appendix C Prompt for Checking Agreement
The prompt for checking agreement is: For each sample, given a set of text sentences, check whether all the following sentences are semantically close to each other or not. You should also consider differences in subtle and nuanced meanings. You should provide an explanation, and then output "yes" if all sentences are semantically close to each other; otherwise, output "no".
Appendix D Prompt for Implication Modeling
[Dialogue context]: {CONTEXT}
[Current turn with emphasis]: {CURRENT}
Appendix E Inter-annotator Agreement for Human Evaluation Score
@@
Appendix F Details of Automatic Evaluation
@@
Appendix G Dataset License
We plan to release the Emphasized-Talk dataset under the MIT license.
Appendix H Details of Human Annotation Process

### v2

@@ Abstract
Guan-Ting Lin, Hung-yi Lee Graduate Institute of Communication Engineering, National Taiwan University Taiwan {f10942104, hungyilee}@ntu.edu.tw
@@ 1 Introduction
Emphasis plays a key role in communication by highlighting important parts of dialogue. Recognizing these emphatic cues is crucial for achieving natural interactions, as they help reinforce intentions, and express nuances essential for understanding conversations. For instance, altering the emphasis in the sentence I never said he stole my bag from he to my can drastically change its meaning. Emphasis can be represented in various ways, such as quotation marks, bold text, italics, capitalization, and underlining.
Large Language Models (LLMs) (Wei et al., 2022; Touvron et al., 2023a; OpenAI, 2023) have revolutionized natural language processing by leveraging vast amounts of unlabeled text data. They can generate and understand natural language accurately, capturing nuances and complex relationships. As a result, LLMs are widely used in applications such as dialogue systems and virtual assistants. Despite their capabilities, the ability of LLMs to grasp the subtle nuances and complex meanings conveyed through emphasis in human dialogue remains unclear.
@@
To address this gap, we introduce a novel benchmark evaluation dataset, Emphasized-Talk, which features real dialogue samples with annotations capturing the implications of emphasis as interpreted by humans. Emphasized-Talk includes the same dialogue context and current sentence but with different words or phrases emphasized. This dataset is designed to test whether LLMs can accurately understand the meaning and intention behind emphasis in dialogue. Note that in this work, we use quotation mark ("") to highlight the emphasis in text sentences. Our study includes both open-sourced and commercial LLMs of varying sizes to provide a comprehensive analysis of their current capabilities and limitations.
Furthermore, we propose an automatic evaluation method that leverages the capabilities of GPT-4 OpenAI (2023) to assess the performance of these models. By comparing various automatic evaluation methods, we find that GPT-4’s evaluations correlate well with human evaluation scores. Our findings indicate that open-source LLMs generally struggle to understand the implications of emphasized sentences. While commercial LLMs perform better, there is still room for improvement in their comprehension of emphasized text.
@@
    We introduce an automated evaluation pipeline using GPT-4 with our benchmark, enabling efficient and consistent assessment of models, which significantly reduces the need for manual testing.
@@ 2 Related works
Human Communication Beyond Textual Information: In addition to text, the extra tags like emotion (Pohl et al., 2017; Xue et al., 2023), emoji (Demszky et al., 2020), sentiment (Lin et al., 2024b), prosody (Ward et al., 2012; Lin et al., 2024a), and non-lexical verbal sounds (Ward, 2006) are critical in understanding and generating human dialogue. These cues provide essential context and pragmatic use to the interaction, often conveying nuances that textual information alone cannot fully capture (Ward, 2004). For instance, emotions and sentiment can indicate the speaker’s attitude and feelings, while speaking style and prosody can reveal additional layers of meaning, such as sarcasm, emphasis, or urgency. This work focuses on emphasis as the non-textual cue in human communication, aiming to understand the implication meaning of beyond text. The quotation mark ("") is used to simulate the expression of emphasis.
Emphasis in Dialogue: In human communication, whether spoken or written, emphasis is a crucial tool for enhancing meaning, indicating emotional states, or highlighting important dialogue segments. In dialogue systems, understanding the subtleties conveyed through emphatic cues is essential for achieving truly natural interactions (Pierrehumbert and Hirschberg, 1990; Wagner and Watson, 2010; Jackson, 2016). Emphasis aids in disambiguating syntactic structures, reinforcing intentions, and expressing nuances that are vital for comprehensive understanding in conversational contexts (Buchanan, 2013). Due to the importance of emphasis, studies on emphasis detection (Arons, 1994; Suni et al., 2017; Talman et al., 2019; Zhang et al., 2018; Vaidya et al., 2022; Morrison et al., 2024), emphatic Text-to-Speech synthesis (Fernandez and Ramabhadran, 2007; Seshadri et al., 2022; Stephenson et al., 2022), and Speech-to-Speech translation (Goldman et al., 2016; de Seyssel et al., 2023) are long-standing research topics to capture emphasis. This work focus on evaluating the LLMs’ ability to understand the high-level meaning and intention of emphasis.
@@ 3.1 Sentence Emphasis Selection
For each dialogue, we select the current sentence to be emphasized based on the following criteria: (1) Sufficient dialogue context: The current turn is selected only if there are more than two preceding dialogue turns, ensuring adequate contextual information. (2) Availability of emphasized targets: Since emphasizing function words like punctuation, articles, and proper nouns rarely affect sentence meaning, we select sentences containing more than four non-functional words and proper nouns for further consideration.
@@ 3.2 Human Annotation with Implication
Selecting which words and phrases to emphasize is non-trivial. In human communication, emphasis placement depends on the dialogue context and the information the speaker wishes to convey. In this work, human annotators choose where to place emphasis and document the implied meaning behind their choices, making the data more pragmatic and reflective of real-world dialogue. On average, the emphasized fragments consist of 1.15 words, indicating that annotators mostly emphasize single words, with occasional emphasis on phrases. Each annotator selects two different words or phrases to emphasize and note the implied meaning. The task template and instructions are shown in Appendix Figure 3.
@@ 3.3 Filtering by Agreement
When multiple annotators select the same word or phrase to emphasize, there are often several annotations of implied meanings. If the implied meanings are significantly different between annotators, indicating a lack of consensus or too many possible implied meanings based on annotators’ backgrounds, we check the semantic similarity among the annotated meanings using GPT-4 (see Appendix B for details). Samples without agreement are filtered out. The percentage of samples that passed the filtering is 79.1%. Overall, we collect 984 dialogue sample pairs. We show some examples in Appendix Table 5.
@@
Table 2: Qualitative example with Emphasizing on different targets in the current turns. 0 and 1 denote the speaker’s identity. For implication rows, the word or phrase within the quotation mark refers to the emphasized target, and the sentence after the colon is the implication. The Predicted implication here is from the Claude 3 Sonnet model.
@@ 4 Evaluation
4.1 Large Language Models to be Evaluated
In our evaluation framework, LLMs only interpret the emphasis as conveyed in the input, but do not detect the emphasis. We use both open-source and closed commercial LLMs for evaluation. Specifically, we evaluate different versions of Llama (Touvron et al., 2023a) (Llama 2 (Touvron et al., 2023b) and Llama 3 (AI@Meta, 2024)) and various model sizes (ranging from 7B to 70B parameters). In addition to Llama, we experiment with Mistral 7B (Jiang et al., 2023). All models used are the instruction-tuned chat versions, not the pre-trained models. For commercial LLMs, we evaluate Claude 3 Sonnet [^2] https://www.anthropic.com/news/claude-3-family and ChatGPT 3.5 (gpt-3.5-turbo-0125) [^3] https://openai.com/index/chatgpt/. For all the LLMs, we provide the following prompt to generate the implied meaning of the emphasized sentence, where {CONTEXT} and {CURRENT} are placeholders that vary for each sample.
[Dialogue context]: {CONTEXT}
[Current turn with emphasis]: {CURRENT}
4.2 Evaluation Metrics
For human evaluation, we recruit human annotators to rate the scoring from 0 to 5, given the rating guideline (the detailed instruction is shown in Appendix Figure 4). Each sample is rated by three annotators, the average score is the Mean Opinion Score (MOS).
However, human evaluation incurs high costs and significant time requirements, requiring skilled evaluators who can understand and analyze the subtle nuances of language and context within the dialogues. Therefore, we propose using the state-of-the-art LLM, GPT-4 (OpenAI, 2023), for automatic evaluation. The automatic evaluation methods are illustrated in Figure 2. We adopt a similar LLM automatic evaluation method as in previous works (Chiang and Lee, 2023a, b), instructing the LLM to predict ratings based on guidelines. This is denoted as auto-gpt4 score. Given potential uncertainties in the LLM’s comprehension of emphasized sentences, we also use a similarity score compared to ground truth annotations, employing both the BERTScore (BERTf1) and GPT-4 (auto-gpt4-gt). BERTScore (Zhang et al., 2020) is a metric for evaluating text generation quality by comparing generated texts to reference texts using BERT embeddings. Unlike traditional metrics like BLEU, which rely on exact matches, BERTScore uses contextual embeddings to capture semantic similarities. As for auto-gpt4-gt, Appendix D describes more details on the prompt template for GPT-4 evaluation.
@@ 5.1 LLMs’ Performance on Implication Modeling
Table reports show the scores of different LLMs. Commercial LLMs generally outperform open-source ones in MOS. The top model is Claude 3 Sonnet with a MOS of 3.73, followed by ChatGPT at 3.59. Among open-source LLMs, Llama 3-70B scores 3.51, close to ChatGPT. The Llama 3 series consistently outperforms the Llama 2 series of similar sizes. Mistral 7B ranks between Llama 2-7B and Llama 3-8B, even surpassing Llama 2-70B. None of the LLMs reach the optimal score of 5, with the best at 3.73, indicating room for improvement in understanding emphasized sentences. More error analyses are shown in Appendix A.
@@ 5.3 Relative Ranking for Automatic Evaluation Methods
The ranking order of LLMs by automatic evaluation scores aligns closely with human scores. Spearman’s rank correlation between human MOS and three automatic methods shows high coefficients: 0.964 for auto-gpt4 and BERTf1, and 0.857 for auto-gpt4-gt. Auto-gpt4 and BERTf1 had only two incorrect model rankings. Despite low Pearson’s and Kendall’s correlations, BERTf1’s rankings are similar to human MOS. Overall, auto-gpt4 achieves both the highest score and rank correlation.
@@ 5.4 Qualitative Example
Table 2 shows examples of the same dialogue context with different emphasized words or phrases. For the emphasized word single, the Claude 3 Sonnet model incorrectly interprets it as referring to a cart for individual use, whereas it actually refers to the number of carts. For the emphasized phrase right now, the model correctly understands it implies immediate use. This demonstrates that while commercial LLMs can infer different meanings based on emphasis, their accuracy depends on the complexity and depth of the context.
@@ Acknowledgement
The authors want to thank Cheng-Han Chiang for his insightful discussion. Guan-Ting Lin is supported by the NTU GICE and the NTU Ph.D. scholarship.
Appendix
@@
Appendix A Analysis on Unsatisfied Model Prediction
We analyze the cases resulting in unsatisfactory ratings (average MOS score below 2). The examples are shown in Table 4. We observe that Claude 3 Sonnet overinterprets the emphasized information in an extremely contrasting way, but this is not relevant to the dialogue context. In contrast, the Llama 2-7B and Mistral-7B models fail to focus on the current turn and the emphasized word, merely summarizing the entire dialogue. These results indicate that small open-sourced LLMs struggle to understand emphasized text and its implications, and sometimes even fail to follow instructions. On the other hand, commercial LLMs can follow instructions but may overinterpret the emphasized text. The ideal interpretation should be reasonable and relevant to the dialogue context.
Appendix B Prompt for Checking Agreement
The prompt for checking agreement is: For each sample, given a set of text sentences, check whether all the following sentences are semantically close to each other or not. You should also consider differences in subtle and nuanced meanings. You should provide an explanation, and then output yes if all sentences are semantically close to each other; otherwise, output no.
Appendix C Inter-annotator Agreement for Human Evaluation Score
@@
Appendix D Details of Automatic Evaluation
@@
Appendix E Dataset License
We release the Emphasized-Talk dataset under the MIT license.
Appendix F Details of Human Annotation Process
@@
Example 1 Input 1. The speaker is expressing genuine happiness and appreciation for the compliment about their country. 2. I am genuinely happy to hear that you love japan. 3. I express my happiness upon hearing your admiration for japan. Output Output: Yes Explanation: All three sentences convey the speaker’s genuine happiness in response to a compliment about Japan. They all express positive feelings related to the admiration of Japan. Example 2 Input 1. You should have expressed your deep feelings for her to me. 2. The relationship is serious and significant. 3. You should have informed me about your romantic feelings for her. Output Output: No Explanation: The second sentence discusses the seriousness of the relationship in general, which differs from the specific recommendation of expressing feelings in the first and third sentences.
Table 5: The examples of filtering by agreement process.
Example of auto-gpt4-gt score Input Dialogue context: What sort of music do you like? Current turn: I love just about “anything” by prince. Model prediction: The speaker has a strong preference for and enjoyment of Prince’s music. Ground truth: The speaker’s affection for prince’s music encompasses a wide range, indicating a broad and inclusive taste. Output Analysis: The ground truth emphasizes the breadth of the speaker’s affection towards Prince’s music, highlighting an inclusive and wide-ranging taste. The model prediction focuses more narrowly, noting the speaker’s strong preference and enjoyment, but doesn’t capture the aspect of ’broad and inclusive taste’. Therefore, while both sentences refer to the speaker’s positive feelings towards Prince’s music, the model prediction lacks the subtlety about the speaker’s taste being wide and inclusive. Score: 3
Table 6: Example input and output for the auto-gpt4-gt score.